# A quasilinear Keller-Segel model with saturated discontinuous advection

### v2

@@ Abstract.
We consider the singular limit of a chemotaxis model of bacterial collective motion recently introduced in [2]. The equation models aggregation-diffusion phenomena with advection that is discontinuous and depends sharply on the gradient of the density itself. The quasi-linearity of the problem poses major challenges in the construction of the solution and complications arise in the proof of regularity. Our method overcomes these obstacle by relying solely on entropy inequalities and the theory of monotone operators. We provide existence, uniqueness and smoothing estimates in any dimensional space.
Key words and phrases:
Degenerate diffusion, parabolic PDEs, Keller-Segel, existence of weak solutions, quasilinear PDEs
M.P.G. is partially supported by NSF Grant DMS-2206677. The authors would like to thank Franca Hoffmann and Inwon Kim for the fruitful discussions.
@@ 1. Introduction and main result
Since the classical experiment of J. Adler on Escherichia coli [1] many experiments have confirmed the formation and propagation of traveling pulses of bacteria. The Keller-Segel model has been widely used as a basis to explain this phenomena in bacterial colonies driven by chemotaxis, see [16, 17]: it describes how the cell density moves towards a nutrient $N$ and a chemical molecule $S$ (chemoatractant) which is generated by the cells themselves. The velocity of this movement is proportional to the gradient of these two quantities. The Keller-Segel model has been the subject of intense mathematical investigation.
In recent works ([4, 19, 14, 10, 5, 7]), a novel class of models has been formulated, originating from a mesoscopic description of the run-and-tumble phenomenon at the individual cell level. Mathematically, these models deviate from the classical Keller-Segel framework in that the cells’ velocity (or flux) is saturated and is proportional to $\frac{\nabla S}{|\nabla S|}$ and $\frac{\nabla N}{|\nabla N|}$. The model, in its more general form, reads as
@@
In [4], the authors present an explicit formula for traveling pulses and establish the existence of a steady state, mirroring the physical occurrence of cluster formation at the bacteria level (e.g., in the absence of nutrients). They also demonstrate the alignment of numerical experiments with actual data. The bounded fluxes, distinct from the flux in Keller-Segel, offer a significant advantage in investigating global well-posedness and, notably, in exploring traveling pulses and long-term asymptotics. Overall, this model has demonstrated greater robustness in replicating the observed traveling bands in experiments.
More recently, Calvez and Hoffmann [2] considered a simplified version of the above model (absence of nutrient and quasi-stationary chemoattractant signal)
$\left\{\begin{array}[]{rcll}\partial_{t}\rho-\Delta\rho+\chi\textrm{div}\,\left(\rho\frac{\nabla S}{|\nabla S|}\right)&=&0,\\ \vspace{2mm}-\Delta S+\alpha S&=&\rho,\end{array}\right.$ (3)
and perform nonlinear stability analysis for the one dimensional case. In particular they show that if the initial data is near, in some norm, to the equilibrium function $\rho_{\infty}=\frac{\chi}{2}e^{-\chi|x-x_{0}|}$, then the corresponding solution converges to $\rho_{\infty}$ exponentially fast. The constant $\alpha\geq 0$ denotes the natural decay of the chemical component.
In this manuscript we deal with the model obtained from (3) by performing the singular limit in $\alpha$. After rescaling $S\mapsto S/\alpha$, the limit $\alpha\rightarrow\infty$ yields the following quasi-linear parabolic problem:
@@
Unlike (3), the advection depends on the gradient of the density, making it a competing term with the diffusion. Note in fact that
$\Delta\rho-\chi\textrm{div}\,\left(\rho\frac{\nabla\rho}{|\nabla\rho|}\right)=\textrm{div}\,\left(\nabla\rho\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)\right),$
which shows a negative diffusion for profiles that have $\chi\rho\geq|\nabla\rho|$. For example, Gaussian functions $e^{-|x|^{2}}$ will make the diffusion coefficient negative around the zero. The system is diffusive as long as $\lvert\nabla\rho\rvert>\chi\rho$, and the steady-state is governed by the equation
@@
for some $c_{1}$, $x_{0}$ determined by the conservation of mass and property of invariance by translation of the equation. Other admissible (i.e. belonging to $L^{1}(\mathbb{R}^{d})$) solutions to (5) exist; for example, for $d=1$ another solution is the “multi-peak profile”
@@
for a suitable choice of $k\geq 0$, $c_{0}\ldots c_{k}>0$, and $x_{0}\ldots x_{k+1}\in\mathbb{R}$, while for $d\geq 2$ another solution is the “factorized profile”
@@
which equals the product of (suitably rescaled) single variable steady states. The proof of existence of solution to (4) is still out of reach. The main difficulty is the control on the diffusion coefficient, which can change along the flow and destroy the parabolic nature of the equation. We believe that if the initial data makes the diffusion coefficient positive, the coefficient will remain non-negative throughout the evolution. This is, however, still a claim. Problem (4) has a more tractable form in one dimension
@@
$\left\{\begin{array}[]{rcll}\partial_{t}\rho-\textrm{div}\,\left[\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\right]&=&0,&\quad\mbox{in }\mathbb{R}^{d}\times(0,\infty),\vspace{2mm}\\ \vspace{2mm}\rho(0)&=&\rho_{in},&\quad\mbox{in }\mathbb{R}^{d},\end{array}\right.$ (9)
where $(\cdot)_{+}$ denotes the positive part. Here we use the convention $f=f_{+}-f_{-}$. Equation (9) is quasilinear. The type of nonlinearity poses major challenges in the construction of solutions. Let us, however, quickly summarize why global well-posedness for (9) is more in-reach. First, observe that the operator
$\left(1-\frac{\chi}{\lvert\nabla\log\rho\rvert}\right)_{+}\nabla\log\rho$
is monotone. More precisely,
$\left[\left(1-\frac{\chi}{\lvert\nabla\log\rho\rvert}\right)_{+}\nabla\log\rho-\left(1-\frac{\chi}{\lvert\nabla\log z\rvert}\right)_{+}\nabla\log z\right]\left(\nabla\log\rho-\nabla\log z\right)\leq 0.$
This monotonicity property, coupled with the theory of monotone operators and a fixed-point argument, proves fundamental in the proof of existence of solutions to (9) equipped with an extra viscosity term $\varepsilon\Delta\rho$. The crux of the analysis will be the vanishing viscosity limit $\varepsilon\to 0$ and the corresponding identification of the limit of the nonlinear term
@@
Before we dwell with analytical details, let us mention that a kinetic formulation and related stability analysis of (3) in one dimensional space has been considered in several works [18, 8, 9, 3]. The singular limit $\alpha\to+\infty$ in this setting is still an open problem.
Our main results are summarized in the following theorem:
@@
Let $\rho_{{\rm in}}\in L^{p}(\mathbb{R}^{d})\cap L^{1}_{2}(\mathbb{R}^{d})$, with $p>d$, $\rho_{{\rm in}}>0$ a.e. in $\mathbb{R}^{d}$ and such that $\int_{\mathbb{R}^{d}}\rho_{{\rm in}}|\log{\rho_{{\rm in}}}|<\infty$. Then, for every $T>0$ there exists an unique solution $\rho:[0,T)\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ to (9) such that
@@
    $\|\nabla\rho_{\varepsilon}\|_{L^{p}((t,T)\times\mathbb{R}^{d})}<C(\rho_{{\rm in}}),\quad\forall\;t>0.$
As mentioned earlier, the most important observation is that the diffusion operator in (9) is degenerate, but monotone. The construction of solution to (9) with extra viscosity and a-priori estimates uniform in the viscosity parameter will follows automatically from the theory of monotone operators. The crux of the analysis is, however, the identification of the limit of the nonlinear term in the vanishing viscosity limit:
A common strategy would involve showing strong convergence of one of the two terms. However, standard regularity techniques, usually based on obtaining estimates for the first derivative through second-order test functions, prove unsuitable here due to the discontinuity of the advection term and the discontinuity of $\nabla\rho$ itself. We overcome this problem using the Boltzmann relative entropy between two viscosity solutions. Let $H[\rho_{\varepsilon},\rho_{\delta}]$ be the nonnegative functional defined as
@@
This functional measures the distance, in $L^{1}$ norm, between the functions $\rho_{\varepsilon}$ and $\rho_{\delta}$. Direct computations show that
@@
$-\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\nabla\log\left(\frac{\rho_{\varepsilon}}{\rho_{\delta}}\right)\right|^{2}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x$
[…]
@@
$-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi\rho_{\delta}}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x\,{\rm d}t\to 0,$
@@
$-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\nabla\log\left(\frac{\rho_{\varepsilon}}{\rho_{\delta}}\right)\right|^{2}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t\to 0.$
Note that $H[\rho_{\varepsilon},\rho_{\delta}](0)=0$ because $\rho_{\varepsilon}$and $\rho_{\delta}$ have same initial data. The two limits above will imply that, as $\varepsilon$ and $\delta$ approach zero,
@@
weakly in $L^{2}$. The identification of this limit is the content of Section 4.1. Almost automatically, the relative entropy argument will also provide uniqueness of the solution for the system without viscosity. In conclusion, our analysis establishes the existence of a unique solution in $L^{p}(t,T,C^{\alpha}(\mathbb{R}^{d}))$ for any $t>0$. We obtain regularity for $\partial_{t}\rho$ in Sobolev spaces with negative coefficients. The question of whether one can obtain more regularity is left for future investigation. Potential insights might be drawn from existing results on the $p$-Laplacian equation [6, 11]. However, the structure of our degenerate set is slightly more complicated than the one of the $p$-Laplacian, namely $|\nabla\rho|=0$. A De Giorgi argument for Hölder regularity might pose challenges, since our diffusion coefficient is degenerate, probably in a set of positive measure, and depends in a nonlinear way on $\nabla\log\rho$.
The proof contains two approximation steps: initially we introduce viscosity and regularize the initial data $f_{in}$ with a sequence of very smooth functions that converge (strongly in $L^{p}$) to $f_{in}$. We establish the existence of solutions to the approximated problem and subsequently remove the viscosity. In the last step we show how the solution obtains the initial data strongly in $L^{p}$ sense.
The following section provides a concise summary of the most relevant a-priori estimates. Section 3 delves into the global well-posedness of the model with viscosity, while Section 4 focuses on the vanishing viscosity limit and establishes uniqueness.
2. A priori estimates and technical results
In this section we give some a priori estimates of the solutions to
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho\varphi dxdt=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\;dxdt$ (11)
[…]
for any $\varphi\in L^{2}(0,T,H^{1}(\mathbb{R}^{d}))$ with initial smooth data $\rho_{0}$. The function $\rho_{0}$ belongs to $W^{1,p}\cap L^{\infty}\cap L^{1}_{2}(\mathbb{R}^{d})$, with $p>d$. The following a priori estimates hold:

### v3

@@ Abstract.
We consider the singular limit of a chemotaxis model of bacterial collective motion recently introduced in [2]. The equation models aggregation-diffusion phenomena with advection that is discontinuous and depends sharply on the gradient of the density itself. The quasi-linearity of the problem poses major challenges in the construction of the solution and complications arise in the proof of regularity. Our method overcomes these obstacle by relying solely on entropy inequalities and the theory of monotone operators. We provide existence, uniqueness and $L^{\infty}$-smoothing estimates in any dimensional space.
M.P.G. is partially supported by NSF Grant DMS-2206677 and DMS-2511625. The authors would like to thank Franca Hoffmann and Inwon Kim for the fruitful discussions.
@@ 1. Introduction and main result
Since the classical experiment of J. Adler on Escherichia coli [1] many experiments have confirmed the formation and propagation of traveling pulses of bacteria. The Keller-Segel model has been widely used as a basis to explain this phenomena in bacterial colonies driven by chemotaxis, see [16, 17]: the model describes how the cell density moves towards a nutrient $N$ and a chemical molecule $S$ (chemoatractant) which is generated by the cells themselves. The velocity of this movement is proportional to the gradient of these two quantities. The Keller-Segel model has been the subject of intense mathematical investigation.
Recent works ([4, 19, 14, 10, 5, 7]) have introduced a novel class of equations, originating from a mesoscopic description of the run-and-tumble phenomenon at the individual cell level. Unlike the classical Keller-Segel framework, these models incorporate a saturated cell velocity (or flux), which is aligned with the gradients of chemical signals, specifically proportional to $\frac{\nabla S}{|\nabla S|}$ and $\frac{\nabla N}{|\nabla N|}$. The general form of this new class of equations is given by:
@@
In [4], the authors derive an explicit expression for traveling pulses and prove the existence of steady states, corresponding to the physical clustering of bacteria (e.g., in nutrient-depleted conditions). Their numerical simulations show strong agreement with experimental data. The use of bounded fluxes, in contrast to Keller-Segel, provides key advantages for proving global well-posedness and for analyzing traveling pulses and long-time dynamics. Overall, this model more accurately captures the experimentally observed traveling bands and offers improvements in describing bacterial movement.
More recently, Calvez and Hoffmann [2] considered a simplified version (absence of nutrient and quasi-stationary chemoattractant signal)
$\left\{\begin{array}[]{rcll}\partial_{t}\rho-\Delta\rho+\chi\textrm{div}\,\left(\rho\frac{\nabla S}{|\nabla S|}\right)&=&0,\\ \vskip 5.69054pt-\Delta S+\alpha S&=&\rho,\end{array}\right$. (3)
and perform nonlinear stability analysis for the one dimensional case. In particular, they show that if the initial data is near, in some norm, to the equilibrium function $\rho_{\infty}=\frac{\chi}{2}e^{-\chi|x-x_{0}|}$, then the corresponding solution converges to $\rho_{\infty}$ exponentially fast. The constant $\alpha\geq 0$ denotes the natural decay of the chemical component and $\chi$ the chemoattractant sensitivity. A kinetic formulation and related stability analysis of (3) in one dimensional space has been considered in [18, 8, 9, 3].
In this manuscript we analyze a model obtained from (3) via the singular limit in $\alpha$. After rescaling $S\mapsto S/\alpha$, the limit $\alpha\rightarrow\infty$ yields the following quasi-linear parabolic problem:
@@
Unlike (3), the advection in (4) depends on the gradient of the density, making it a competing term with the diffusion. Note in fact that
$\Delta\rho-\chi\textrm{div}\,\left(\rho\frac{\nabla\rho}{|\nabla\rho|}\right)=\textrm{div}\,\left(\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)\nabla\rho\right),$
which shows a negative diffusion for profiles that have $\chi\rho\geq|\nabla\rho|$. For example, Gaussian functions $e^{-|x|^{2}}$ will make the diffusion coefficient negative around $|x|=0$. Problem (4) is diffusive as long as $\lvert\nabla\rho\rvert>\chi\rho$, and the steady-state is governed by the equation
@@
for some $c$ and $x_{0}$ determined by the conservation of mass and property of invariance by translation of the equation. Other admissible (i.e. belonging to $L^{1}(\mathbb{R}^{d})$) solutions to (5) exist; for example, for $d=1$ another solution is the “multi-peak profile”
@@
for a suitable choice of $k\geq 0$, $c_{0}\ldots c_{k}>0$, and $x_{0}\ldots x_{k+1}\in\mathbb{R}$, while for $d\geq 2$ other possible solution to (5) is the “factorized profile”
@@
[…]
@@
$\left\{\begin{array}[]{rcll}\partial_{t}\rho-\textrm{div}\,\left(\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\right)&=&0,&\quad\mbox{in }\mathbb{R}^{d}\times(0,\infty),\vskip 5.69054pt\\ \vskip 5.69054pt\rho(0)&=&\rho_{in},&\quad\mbox{in }\mathbb{R}^{d},\end{array}\right$. (9)
where $(\cdot)_{+}$ denotes $\max\{\cdot,0\}$. Equation (9) is a quasilinear differential equations. While this nonlinearity still poses major challenges in the construction of solutions, let’s explain why (9) is more within-reach than (4). The most important property of (9) is that the operator
$\left(1-\frac{\chi}{|w|}\right)_{+}w,\quad w\in\mathbb{R}^{d},$
is monotone (for any constant $\chi>0$). More precisely,
$\left[\left(1-\frac{\chi}{|w|}\right)_{+}w-\left(1-\frac{\chi}{|z|}\right)_{+}z\right]\cdot(w-z)\;{{\geq}}\;0,\quad\quad\forall\;w,z\in\mathbb{R}^{d}.$ (10)
Note that the monotonicity is lost if we remove the positive part. This monotonicity will prove fundamental in two steps: (i) in the construction of solutions to a viscous version of (9), and, most importantly, (ii) in the vanishing viscosity limit. The crux of the analysis will be the vanishing viscosity limit $\varepsilon\to 0$ and the corresponding identification of the limit of the nonlinear term
@@
We will expand on this point after we summarize our main result:
@@
Let $\rho_{{\rm in}}\in L^{p}(\mathbb{R}^{d})\cap L^{1}_{2}(\mathbb{R}^{d})$, $d\geq 1$, with $p>d$, $\rho_{{\rm in}}>0$ a.e. in $\mathbb{R}^{d}$ and $\int_{\mathbb{R}^{d}}\rho_{{\rm in}}|\log{\rho_{{\rm in}}}|<\infty$. For every $T>0$ there exists an unique solution $\rho:[0,T)\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ to (9) such that
@@
    $\|\nabla\rho\|_{L^{p}((t,T)\times\mathbb{R}^{d})}<C(\rho_{{\rm in}}),\quad\forall\;t>0.$
A common strategy for proving existence in nonlinear problems involves initially linearizing the equation and then applying fixed-point or contraction mapping arguments. If we were to follow this approach, we would begin by linearizing the diffusion term in (11) as
$\left(1-\chi\frac{z}{\lvert\nabla z\rvert}\right)_{+}\nabla\rho.$
However, because this linearization includes the term $|\nabla z|$, the functional framework for any fixed-point or contraction method would necessarily involve the Sobolev space $H^{1}$. The argument would also require regularity estimates in a higher space, typically $H^{2}$. This poses a problem, as we do not expect solutions to be in $H^{2}$. In fact, at points where $|\nabla\rho|=0$, we anticipate a loss of smoothness. Some steady states of type (6) are no more regular than Lipschitz continuous. Therefore, this standard linearization route is not suitable here.
Instead, our approach relies on the monotonicity property (10), through a different form of linearization. Specifically, we employ a semi-linearization of the diffusion term in (11):
$\left(1-\chi\frac{z}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho.$
This operator in monotone in the variable $\nabla\rho$, see (10). Existence of solutions to the semi-linearized problem follows then from the theory of monotone operators, after adding some viscosity terms $\varepsilon(\Delta\rho-\rho)$ to secure uniform ellipticity. In this framework, fixed-point arguments will be carried out in $L^{2}$ and require $H^{1}$-type of estimates for $\rho$.
The next crucial step involves taking the vanishing viscosity limit, particularly identifying the limit of the nonlinear flux:
$\lim_{\varepsilon\to 0}\left(1-\frac{\chi}{\lvert\nabla\log\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}.$
@@
To overcome this, we use the Boltzmann entropy between two viscous solutions. The Boltzmann relative entropy between two solutions $\rho_{\varepsilon}$ and $\rho_{\delta}$ is defined as
$H[\rho_{\varepsilon},\rho_{\delta}]=\int_{\mathbb{R}^{d}}\rho_{\delta}\left[\frac{\rho_{\varepsilon}}{\rho_{\delta}}\log{\frac{\rho_{\varepsilon}}{\rho_{\delta}}}-\frac{\rho_{\varepsilon}}{\rho_{\delta}}+1\right]\;dx.$
This functional bounds the in $L^{1}$-distance between $\rho_{\varepsilon}$ and $\rho_{\delta}$. Direct computations show that
@@
$-\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\nabla\log\left(\frac{\rho_{\varepsilon}}{\rho_{\delta}}\right)\right|^{2}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\hskip-7.11317pt+\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x$
$+\textrm{error terms},$
@@
$\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi\rho_{\delta}}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x\,{\rm d}t\to 0,$
@@
$\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\nabla\log\left(\frac{\rho_{\varepsilon}}{\rho_{\delta}}\right)\right|^{2}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\hskip-7.11317pt+\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t\to 0.$
Note that $H[\rho_{\varepsilon},\rho_{\delta}](0)=0$ because $\rho_{\varepsilon}$and $\rho_{\delta}$ have same initial data. The main point here is that the two limits above will imply that, as $\varepsilon$ and $\delta$ approach zero,
@@
weakly in $L^{2}$. The identification of this limit is the content of Section 4.1 and Section 4.2. Almost automatically, the relative entropy argument will also provide uniqueness of the solution for the system without viscosity.
The question of whether one can obtain more regularity is left for future investigation. Based on steady-state solutions of the form (6), the maximal regularity one should expect is Lipschitz continuity. The solution that we obtain is bounded and belongs to $L^{p}(t,T,C^{0,\alpha}(\mathbb{R}^{d}))$ for any $t>0$. Potential insights might be drawn from existing results on the $p$-Laplacian equation [6, 11]. However, the structure of our degenerate set is slightly more complicated than the one of the $p$-Laplacian. A De Giorgi argument for Hölder regularity might pose challenges, since our diffusion coefficient is degenerate, probably in a set of positive measure, and depends in a nonlinear way on $\nabla\log\rho$.
Another very interesting open question is the long time dynamics for (9). While steady states of the form (6) are natural candidates, the set of possible profiles also include all functions satisfying $|\nabla\log\rho_{\infty}|\leq\chi$.
Organization of the paper: The rest of the paper is organized as follows: in the next section we present the relevant a-priori estimates. The proof of Theorem 1 is in Section 3 and Section 4: in Section 3 we provide the proof of existence of solutions to the model equipped with extra viscosity. In Section 4 we perform the vanishing viscosity limit and show uniqueness of solution.
We will use the decomposition
$f=f_{+}-f_{-},\quad f_{+}:=\max\{f,0\},\quad f_{-}:=-\min\{f,0\}.$
2. A priori estimates
This section is devoted to a-priori estimates of solutions to (11). We combine, without creating confusion, estimates for (11) and for its viscous version, namely
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho_{\varepsilon}\varphi dxdt=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\;dxdt$ (12)
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\varepsilon}\nabla\varphi\;dxdt-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\;dxdt.$
Notice that (12) coincides with (11) if $\varepsilon=0$. All our estimates in this section hold for any $\varepsilon\geq 0$.